# Variance-Reducing Couplings for Random Features: Perspectives from Optimal Transport

### v1

@@ 3.2 Pushing further with numerical OT solvers
Multi-marginal OT. The insights of Sec. 3.1 have equipped us with pairwise norm-coupled RFs: a computationally efficient coupling that guarantees strictly lower kernel estimator variance. We obtained it by solving the variance reduction OT problem exactly in $m=2$, then combining $\lfloor\frac{d}{2}\rfloor$ independent copies of such pairs to get the ensemble. A natural question is whether we can do better by inducing dependencies between the all the $d$ frequencies’ norms. Solving this multi-marginal optimal transport problem analytically is a challenging open problem.
Copulas as numerical OT solvers. Whilst a full analytic solution to the multi-marginal OT variance-reduction problem is (for now) out of reach, we can make progress using a numerical OT solver. Our basic strategy is to restrict $\Lambda_{m}(\chi_{d})$, the full set of joint distributions over $m$ random variables with $\chi_{d}$ marginals, to a tractable subset amongst which we can efficiently optimise and sample. One such subset is provided by Gaussian copulas (Nelsen, 2006; Haugh, 2016): joint distributions obtained by taking a multivariate Gaussian and pushing each of its coordinates forward first with the Gaussian CDF $F_{\mathcal{N}}$, and then the $\chi_{d}$ inverse CDF $F_{\chi_{d}}^{-1}$. If the diagonal terms of the underlying Gaussian covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{m\times m}$ are equal to $1$ (i.e. it is a correlation matrix), this has the prescribed marginals so unbiasedness is baked in. Meanwhile, correlations between the random variables are controlled by the off-diagonal entries of $\mathbf{\Sigma}$. This family parameterises a very broad (though not universal) set of couplings, including pairwise norm-coupled RFs (Def. 3.3).
In App. A.5 we show that it is possible to use gradient descent with the reparameterisation trick to learn the optimal copula covariance matrix $\mathbf{\Sigma}$, approximately solving the multi-marginal OT problem (Chi et al., 2019). We do this by minimising the kernel approximation error on training data, exploiting the fact that all operations to construct the features are differentiable. Remarkably, optimising the copula recovers our pairwise norm-coupling scheme (or other equally well-performing variants). This suggests that PNC may in fact exactly solve the multi-marginal OT problem in Eq. 6, a tantalising possibility which we aim to explore in future work.
@@ 3.3 Experiments for norm-coupled RFs
Table 1: Performance of RFFs and RLFs on kernel estimation with UCI datasets with different coupling schemes, taking $d$ random frequencies for RFFs and $2d$ for RLFs (see main text for details). We show RMSEs to the ground truth kernel values, normalised such that the RMSE of the i.i.d. estimator is equal to one. Lower is better. Error bars are standard errors on the reported RMSEs.
To verify the efficacy of pairwise norm-coupled RFs (Def. 3.3), we now compute kernel estimates with RFFs and RLFs for real-world UCI datasets. We choose the kernel lengthscale parameters (equivalently, the data normalisation) based on a training set, either by training a GP (RFFs) or selecting reasonable values for downstream Performer applications (RLFs) (Choromanski et al., 2020). We then compute the kernel approximation RMSE on a test set. Full details are in App. B.1.
@@
Results. Table 1 shows the results. For RFFs, we take $m=d$ orthogonal frequencies. For RLFs, we also include their antiparallel directions, giving $m=2d$ frequencies. For each dataset, the RMSEs are normalised by the result with i.i.d. features. As a further baseline, we include RFs constructed using Halton sequences (Dick et al., 2013; Yang et al., 2014a), a fixed, off-the-shelf QMC scheme that sometimes provides gains but is clearly suboptimal. The third row shows orthogonal frequencies with independent norms (Yu et al., 2016). When we also couple the frequencies’ norms using our OT-driven PNC scheme (plus antithetic sampling for RLFs, due to Thm 3.5), we access even lower estimator variance at no extra computational cost. Note that the $z^{2}/d\to 0$ condition for RFFs is found to be nonrestrictive, and coincides with the $d\to\infty$ limit where standard QMC may fail.
Downstream tasks. We have achieved our objective of variance reduction, a popular goal in the literature (Yu et al., 2016; Lyu, 2017). One might therefore expect PNC RFs to improve downstream performance. We do not observe this (see App. B.2): a surprising observation we return to in Sec. 5.
@@ 4 Graph random features
We now shift our attention from $\mathbb{R}^{d}$ to the discrete domain. Consider an undirected graph $\mathcal{G}(\mathcal{N},\mathcal{E})$ where $\mathcal{N}\coloneqq\{v_{1},...,v_{N}\}$ is the set of nodes and $\mathcal{E}$ is the set of edges, with $(v_{i},v_{j})\in\mathcal{E}$ if and only if there exists an edge between $v_{i}$ and $v_{j}$ in $\mathcal{G}$. Graph node kernels $k:\mathcal{N}\times\mathcal{N}\to\mathbb{R}$ are positive definite, symmetric functions defined on pairs of nodes of $\mathcal{G}$, reflecting some notion of their ‘closeness’ via the graph edges and weights. $k$ captures the structure of $\mathcal{G}$, letting practitioners repurpose popular kernelised learning algorithms – including their theoretical guarantees and empirical success – to the discrete domain (Smola and Kondor, 2003b, a). Examples include the diffusion, regularised Laplacian, cosine and random walk kernels, all of which are typically considered as functions of the graph Laplacian matrix (Kondor and Lafferty, 2002). We give a short introduction in App. C.1.
@@
Graph random features. As in the Euclidean setting, graph-based kernel methods scale notoriously poorly on account of the $\mathcal{O}(N^{3})$ time complexity of inverting the Gram matrix $\mathbf{K}\coloneqq[k(v_{i},v_{j})]_{i,j=1}^{N}$. In fact, even computing $\mathbf{K}$ often incurs a cubic cost since it involves multiplying large adjacency matrices. This has motivated research dedicated to improving efficiency by approximating $\mathbf{K}$, including the recently-introduced class of graph random features (GRFs) (Choromanski, 2023; Reid et al., 2024b). These are finite-dimensional, explicitly manifested vectors $\{\phi_{\textrm{GRF}}(v_{i})\}_{i=1}^{N}\subset\mathbb{R}^{N}$ whose Euclidean dot product is equal to the graph kernel evaluation in expectation,
$\mathbf{K}_{ij}=\mathbb{E}\left[\phi_{\textrm{GRF}}(v_{i})^{\top}\phi_{\textrm{GRF}}(v_{j})\right].$ (9)
GRFs enable practitioners to compute a sparse, unbiased estimate of the graph kernel (or its inverse) in subquadratic time. They can be constructed for arbitrary functions of a weighted adjacency matrix.
Coupled random walks. Whereas RFFs and RLFs sample random vectors in $\mathbb{R}^{d}$, GRFs sample simple random walks. These are sequences of graph nodes $(v_{i})_{i=1}^{l}$ where consecutive nodes are connected by an edge, $(v_{i},v_{i+1})\in\mathcal{E}$. At every timestep, the walker chooses one of its neighbours uniformly and at random. The length of the walk $l$ may itself be a random variable. For GRFs, it is drawn from a geometric distribution $l\sim\textrm{G}(p),p\in(0,1)$, which corresponds to the walker terminating with probability $p$ at every timestep.
@@
Random walks are usually taken to be independent, but this can lead to slow mixing times, poor efficiency and high kernel estimator variance (Alon et al., 2007; Zhou et al., 2015). A natural question is: can we use insights from OT to couple graph random walks to improve the convergence of GRFs?
@@
The simplest approach to induce dependencies between random walks is to couple their lengths. [^1] Walker directions can also be coupled for additional variance reduction (Reid et al., 2024a), but our work will not consider (and can be easily combined with) this complementary approach. For GRFs, the constraint for unbiasedness is that the marginal distribution of each random variable $l$ must remain geometric. Reid et al. (2024c) recently proposed a simple, heuristically-motivated algorithm to achieve this called antithetic termination, which directly anticorrelates the termination events between walkers at every timestep (see App. C.2). They provide asymptotic theoretical guarantees, but only for particular choices of graph kernel. In the following section, we will see that an OT-based approach that optimises a coupling using real data performs substantially better.
@@ 4.1 Approximating and solving the OT problem for GRFs
Here, we present our approach for formulating and approximately solving the variance reduction OT problem with GRFs. It works by mapping to a corresponding bipartite matching problem which we can solve efficiently.
Constructing GRFs from random walks. GRFs are constructed differently to RFFs and RLFs. To obtain $\phi_{\textrm{GRF}}(v_{i})\in\mathbb{R}^{N}$ (where $v_{i}\in\mathcal{N}$), we sample $m$ random walks $\{\bm{\omega}_{k}^{(i)}\}_{k=1}^{m}$ out of node $v_{i}$ and compute the arithmetic mean of their corresponding ‘projections’,
@@
Here, the projection function $\psi(\cdot):\Omega\to\mathbb{R}^{N}$ maps from the set of graph random walks $\Omega\coloneqq\left\{(v_{i})_{i=1}^{l}\,|\,v_{i}\in\mathcal{N},(v_{i},v_{i+1})\in\mathcal{E},l\in\mathbb{N}\right\}$ to a sparse $N$-dimensional feature vector satisfying $k(v_{i},v_{j})=\mathbb{E}_{\bm{\omega}^{(i)},\bm{\omega}^{(j)}}[\psi(\bm{\omega}^{(i)})^{\top}\psi(\bm{\omega}^{(j)})]$. $\psi(\cdot)$ depends on the particular kernel being approximated. We refer the reader to the works of Choromanski (2023) and Reid et al. (2024b) for an introduction to GRFs, but also provide background in App. C.1 for completeness. For our purposes, it will be sufficient to note that $\psi(\cdot)$ is a complicated function; it is difficult to reason about analytically but straightforward to compute for a particular walk. Moreover, its input walks are discrete random variables so $\psi(\cdot)$ is not differentiable with respect to their lengths $l_{k}^{(i)}$ (where $k=1,...,m$ and $v_{i}\in\mathcal{N}$ is the walker’s start node) – a property that will preclude optimisation using copulas (Sec. 3.2).
A pair of walkers. Let us initially consider just $m=2$ walkers. The kernel estimator $\widehat{k}(v_{i},v_{j})$ is
@@
which is unbiased provided (i) the marginal distribution of each $\bm{\omega}_{1,2}^{(i,j)}$ is a simple random walk with geometrically distributed length (hereafter denoted $\eta_{\textrm{G}}$) and (ii) walks from node $v_{i}$ are independent from walks from node $v_{j}$. The variance of the estimator depends on the quantity
$\mathbb{E}\left(\widehat{k}(v_{i},v_{j})^{2}\right)=\frac{1}{4}\mathbb{E}_{\bm{\omega}_{1,2}^{(i,j)}}\left[\left(\psi(\bm{\omega}_{1}^{(i)})+\psi(\bm{\omega}_{2}^{(i)})\right)^{\top}\left(\psi(\bm{\omega}_{1}^{(j)})+\psi(\bm{\omega}_{2}^{(j)})\right)\right]^{2},$ (12)
@@
where the expectation is taken over both the directions and lengths of the random walks. Suppose the directions remain independent, but the lengths $l_{1}^{(i)}$ and $l_{2}^{(i)}$ (and likewise $l_{1}^{(j)}$ and $l_{2}^{(j)}$) are to be coupled for variance reduction. The OT problem can then be written
$\mu^{*}=\arg\inf_{\mu\in\Lambda_{2}(\eta_{\textrm{G}})}\mathbb{E}_{(l_{1}^{(i)},l_{2}^{(i)})\sim\mu,(l_{1}^{(j)},l_{2}^{(j)})\sim\mu}\left(\mathbb{E}_{\textrm{dirs}}\left[\left(\psi(\bm{\omega}_{1}^{(i)})+\psi(\bm{\omega}_{2}^{(i)})\right)^{\top}\left(\psi(\bm{\omega}_{1}^{(j)})+\psi(\bm{\omega}_{2}^{(j)})\right)\right]^{2}\right),$ (13)
where $\mathbb{E}_{\textrm{dirs}}$ denotes the expectation over the walkers’ directions, i.e. which neighbour they choose at each timestep.

### v2

@@ 3.2 Pushing further with numerical OT solvers
Multi-marginal OT. In Sec. 3.1 we proposed PNC RFs: a computationally efficient coupling that is guaranteed to reduce variance for any $m$. We obtained it by solving the variance reduction OT problem exactly in $m=2$, then combining $\lfloor\frac{d}{2}\rfloor$ independent copies to get the ensemble. Can we do better by inducing dependencies between the all the $m$ frequencies’ norms? Solving this multi-marginal OT problem analytically is a tough open problem.
Copulas as numerical OT solvers. Whilst an analytic solution to the multi-marginal OT variance reduction problem is (for now) out of reach, we can make progress using a numerical OT solver. Our strategy is to restrict $\Lambda_{m}(\chi_{d})$, the full set of joint distributions over $m$ random variables with $\chi_{d}$ marginals, to a tractable subset amongst which we can efficiently optimise and sample. One such subset is provided by Gaussian copulas (Nelsen, 2006; Haugh, 2016): joint distributions obtained by taking a multivariate Gaussian and pushing each of its coordinates forward first with the Gaussian CDF $F_{\mathcal{N}}$, and then the $\chi_{d}$ inverse CDF $F_{\chi_{d}}^{-1}$. Fig. 1 gives a visual overview for $d=2$. If the diagonal terms of the underlying Gaussian covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{m\times m}$ are equal to $1$ (i.e. it is a correlation matrix), this has the prescribed marginals so unbiasedness is baked in. Meanwhile, correlations between the random variables are controlled by the off-diagonal entries of $\mathbf{\Sigma}$. This parameterises a broad set of couplings, including PNC (Def. 3.3). In App. A.5 we demonstrate that it is possible to use gradient descent with the reparameterisation trick to learn the optimal copula covariance matrix $\mathbf{\Sigma}$, approximately solving the multi-marginal OT problem. We do this by minimising the kernel approximation error on training data, exploiting the fact that all operations to construct the features are differentiable. In doing so, we optimise the RF coupling. Remarkably, this data-dependent optimisation does not to find couplings much better than PNC: see the training curves in Fig. 5. This suggests that our scheme may already be close optimal for $m\neq 2$. Intuitively, one cannot simultaneously anticorrelate too many random variables, so strong pairwise couplings already perform very well. Whilst copulas have previously been used as numerical OT solvers (Chi et al., 2019), this is (to our knowledge) their first application to learning a Monte Carlo coupling.
[…]
copula_schematic
Figure 1: Copula schematic for $d=2$. Random variables are drawn from a Gaussian distribution with correlation matrix $\mathbf{\Sigma}$. They are pushed forward using $F_{\mathcal{N}}$ then $F_{\chi_{d}}^{-1}$ to obtain coupled variables with marginal $\chi_{d}$ distributions. $\mathbf{\Sigma}$ is learned using gradient-based optimisation, approximately solving the multi-marginal OT problem in Eq. 5.
@@ 3.3 Experiments for norm-coupled RFs
To test PNC RFs (Def. 3.3), we now compute kernel estimates with RFFs and RLFs for UCI datasets. We choose the kernel lengthscale parameters based on a training set, by training a GP (RFFs) or selecting reasonable values for Performers (RLFs) (Choromanski et al., 2020). We then compute the kernel approximation RMSE on a test set. Full details are in App. B.1.
@@
Table 1: Performance of RFFs and RLFs on kernel estimation with UCI datasets with different coupling schemes, taking $d$ random frequencies for RFFs and $2d$ for RLFs (see main text for details). We show RMSEs to the ground truth kernel values, normalised such that the RMSE of the i.i.d. estimator is equal to one. Lower is better. Error bars are standard errors on the reported RMSEs. Our couplings consistently give the lowest variance.
@@
Results for variance reduction. Table 1 shows the results. For RFFs, we take $m=d$ orthogonal frequencies. For RLFs, we also include their antiparallel directions, giving $m=2d$ frequencies. For each dataset, the RMSEs are normalised by the result with i.i.d. features. As a further baseline, we include RFs constructed using Halton sequences (Dick et al., 2013; Yang et al., 2014a), a fixed, off-the-shelf QMC scheme that can provide small gains but is clearly suboptimal. The third row shows orthogonal frequencies with independent norms (Yu et al., 2016). When we also couple the frequencies’ norms using our PNC scheme (plus antithetic sampling for RLFs, due to Thm 3.5), we access even lower estimator variance at no extra computational cost. Note that the small $z$ condition for RFFs is found to be nonrestrictive in practice.
Downstream tasks. We have achieved our objective of variance reduction, a popular and intensely studied goal in the literature (Yu et al., 2016; Rowland et al., 2018; Reid et al., 2023; Likhosherstov et al., 2022; Yang et al., 2014a; Le et al., 2013; Bojarski et al., 2017; Choromanski et al., 2017; Lyu, 2017; Shen et al., 2017; Dao et al., 2017; Munkhoeva et al., 2018). It is conventionally understood that PNC RFs should therefore improve downstream performance in applications. Surprisingly, when we run exhaustive Gaussian process experiments in App. B.2, we do not observe such a gain.
The reason for this counterintuitive behaviour is as follows. When optimising a coupling, we minimise the variance of pointwise kernel estimates $\{k(\bm{x}_{i},\bm{x}_{j})\}_{i,j=1}^{N}$. However, functions like the predictive mean and KL divergence are highly nonlinear in these estimates. For example, they may involve inverting a Gram matrix. Downstream quantities therefore depend on the joint distribution of the kernel estimates, which are modified nontrivially by the coupling. Variance reduction alone cannot guarantee an improvement.
Performers. As a concrete example, consider estimating attention, $a_{ij}\coloneqq k(\bm{x}_{i},\bm{x}_{j})/\sum_{l=1}^{N}k(\bm{x}_{i},\bm{x}_{l})$ using random Laplace features (Choromanski et al., 2020). This normalises the kernel evaluation between the $i$ and $j$ tokens by the sum with all the other tokens. Taylor expanding, if the kernel estimators have equal means $\mu$, the average mean square error ${\textrm{MSE}(\widehat{a}_{i})}\coloneqq\frac{1}{N}\sum_{j=1}^{N}\textrm{MSE}(\widehat{a}_{ij})$ obeys
${\textrm{MSE}(\widehat{a}_{i})}=\frac{1}{N^{2}\mu^{2}}\left(\frac{1}{N}\sum_{j=1}^{N}\textrm{Var}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))-\frac{1}{N^{2}}\sum_{j_{1},j_{2}=1}^{N}\textrm{Cov}(\widehat{k}(\bm{x}_{i},\bm{x}_{j_{1}}),\widehat{k}(\bm{x}_{i},\bm{x}_{j_{2}}))\right)+\mathcal{O}(\frac{1}{N^{3}}).$ (8)
Table 2: Performer test accuracies on ImageNet with different coupling schemes. Counterintuitively, maximising the pointwise kernel estimator variance by positively correlating feature norms boosts performance – a different OT problem to the naive, obvious choice.
Laplace Features Test acc.
Orthogonal $0.625$$\pm 0.003$
Orthogonal + PNC $0.620$$\pm 0.003$
Orthogonal + PM $\mathbf{0.633}$$\pm 0.003$
By coupling the frequency norms, PNC reduces $\textrm{Var}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))$ as intended. However, it also reduces the covariance $\textrm{Cov}(\widehat{k}(\bm{x}_{i},\bm{x}_{j_{1}}),\widehat{k}(\bm{x}_{i},\bm{x}_{j_{2}}))$, so ${\textrm{MSE}(\widehat{a}_{i})}$ does not actually substantially improve overall. In stark contrast, if we instead take the positive monotone (PM) coupling where $\{\omega_{i}\}_{i=1}^{m}$ are all equal almost surely, then $\textrm{Var}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))$ is maximised (see App. B.3). But these strong, positive correlations also increase $\textrm{Cov}(\widehat{k}(\bm{x}_{i},\bm{x}_{j_{1}}),\widehat{k}(\bm{x}_{i},\bm{x}_{j_{2}}))$ by an even greater amount. Hence, we find that ${\textrm{MSE}(\widehat{a}_{i})}$ falls (see Fig. 7). This is surprising: maximising the pointwise kernel estimator variance by solving the OT problem with the ‘wrong’ sign on the cost function reduces the MSE of the attention scores after normalisation. In fact, the improvement is so big that it increases the average test accuracy of Performers trained on ImageNet (Deng et al., 2009) by +0.8%, whereas PNC makes no statistically significant difference. See Table 2. This demonstrates the limitations of simple variance reduction and invites a more careful treatment, considering the downstream quantities of interest. App. B.3 gives further discussion and transformer training details.
@@ 4 Graph random features
We now shift our attention from $\mathbb{R}^{d}$ to the discrete domain. Consider an undirected graph $\mathcal{G}(\mathcal{N},\mathcal{E})$ where $\mathcal{N}\coloneqq\{v_{1},...,v_{N}\}$ is the set of nodes and $\mathcal{E}$ is the set of edges, with $(v_{i},v_{j})\in\mathcal{E}$ if and only if there exists an edge between $v_{i}$ and $v_{j}$ in $\mathcal{G}$. Graph node kernels $k:\mathcal{N}\times\mathcal{N}\to\mathbb{R}$ are positive definite, symmetric functions defined on pairs of nodes of $\mathcal{G}$, reflecting some notion of their ‘closeness’ via the graph edges and weights. $k$ captures the structure of $\mathcal{G}$, letting practitioners repurpose popular kernelised learning algorithms to the discrete domain (Smola and Kondor, 2003b; a). Examples include the diffusion, regularised Laplacian, cosine and random walk kernels, all of which are typically considered as functions of the graph Laplacian matrix (Kondor and Lafferty, 2002). We give a short introduction in App. C.1.
@@
Graph random features. As in the Euclidean setting, graph kernel methods scale poorly due to the $\mathcal{O}(N^{3})$ time complexity of inverting the Gram matrix $\mathbf{K}\coloneqq[k(v_{i},v_{j})]_{i,j=1}^{N}$. In fact, even computing $\mathbf{K}$ often incurs a cubic cost since it involves multiplying large adjacency matrices. Research has been dedicated to improving efficiency by approximating $\mathbf{K}$, including graph random features (GRFs) (Choromanski, 2023; Reid et al., 2024b). These are sparse vectors $\{\phi_{\textrm{GRF}}(v_{i})\}_{i=1}^{N}\subset\mathbb{R}^{N}$ that satisfy $\mathbf{K}_{ij}=\mathbb{E}\left[\phi_{\textrm{GRF}}(v_{i})^{\top}\phi_{\textrm{GRF}}(v_{j})\right]$, so their dot product is equal to the true graph kernel in expectation. GRFs are constructed in subquadratic time.
@@
Coupled random walks. For GRFs, the ‘frequencies’ $\{\bm{\omega}_{i}\}_{i=1}^{m}$ are simple random walks: sequences of graph nodes $(v_{i})_{i=1}^{l}$ with $(v_{i},v_{i+1})\in\mathcal{E}$. At every timestep, the walker chooses one of its neighbours uniformly and at random. The length of the walk $l$ is also a random variable, drawn from a geometric distribution $l\sim\textrm{G}(p),p\in(0,1)$. In other words, the walker terminates with probability $p$ at every timestep. Random walks are usually taken to be independent, but this can lead to slow mixing times, poor efficiency and high kernel estimator variance (Alon et al., 2007; Zhou et al., 2015). A natural question is: can we couple graph random walks to improve the convergence of GRFs?
@@
Sec. 3 couples Gaussian vector norms. A simple, analogous approach to couple random walks is via their lengths. For GRFs, the constraint for unbiasedness is that the marginal distribution of each variable $l$ must remain geometric. Reid et al. (2024c) proposed a simple, ad-hoc algorithm to achieve this called antithetic termination, which directly anticorrelates the walkers’ termination events at every timestep (see App. C.2). They provide asymptotic theoretical guarantees, but only for particular choices of graph kernel. In the next section, we will see that our novel method of optimising a coupling with data performs much better.
[…]
grfs_master_schematic
Figure 2: Schematic overview of $\sigma$-coupled GRFs with $\sigma=24351$. Left: permutation density with uniform marginals. Centre: bipartite matching between quantiles of geometric distributions over walk lengths. Right: ensemble of graph random walks with lengths to be coupled.
@@ 4.1 Approximating and solving the OT problem for GRFs
Here, we present our novel approach for formulating and approximately solving the variance reduction OT problem with GRFs. It works by mapping to a corresponding bipartite matching problem which we can solve efficiently with linear programming techniques. Fig. 2 gives a visual overview.
Constructing GRFs from random walks. To obtain $\phi_{\textrm{GRF}}(v_{i})\in\mathbb{R}^{N}$, one samples $m$ random walks $\{\bm{\omega}_{k}^{(i)}\}_{k=1}^{m}$ out of node $v_{i}\in\mathcal{N}$ and averages their ‘projections’,
@@
The projection function $\psi(\cdot):\Omega\to\mathbb{R}^{N}$ maps from the set of graph random walks $\Omega\coloneqq\left\{(v_{i})_{i=1}^{l}\,|\,v_{i}\in\mathcal{N},(v_{i},v_{i+1})\in\mathcal{E},l\in\mathbb{N}\right\}$ to a sparse $N$-dimensional feature vector satisfying $k(v_{i},v_{j})=\mathbb{E}_{\bm{\omega}^{(i)},\bm{\omega}^{(j)}}[\psi(\bm{\omega}^{(i)})^{\top}\psi(\bm{\omega}^{(j)})]$. $\psi(\cdot)$ depends on the particular kernel being approximated. We direct the reader to the work of Reid et al. (2024b) for an introduction to GRFs, and also provide background in App. C.1. $\psi(\cdot)$ is a complicated function; it is difficult to reason about analytically but straightforward to compute for a particular walk. Moreover, its input walks are discrete random variables so $\psi(\cdot)$ is not differentiable with respect to the lengths $l_{k}^{(i)}$ (where $k=1,...,m$ and $v_{i}\in\mathcal{N}$ is the walker’s start node). This precludes straightforward gradient-based optimisation (Sec. 3.2).
A pair of walkers. Initially consider just $m=2$ walkers. The kernel estimator $\widehat{k}(v_{i},v_{j})$ is
@@
which is unbiased provided (i) the marginal distribution of each $\bm{\omega}_{1,2}^{(i,j)}$ is a simple random walk with geometrically distributed length (hereafter denoted $\eta_{\textrm{G}}$) and (ii) walks from node $v_{i}$ are independent from walks from node $v_{j}$. The variance of the estimator depends on
$\mathbb{E}\left(\widehat{k}(v_{i},v_{j})^{2}\right)=\frac{1}{16}\mathbb{E}_{\bm{\omega}_{1,2}^{(i,j)}}\left(\left[\left(\psi(\bm{\omega}_{1}^{(i)})+\psi(\bm{\omega}_{2}^{(i)})\right)^{\top}\left(\psi(\bm{\omega}_{1}^{(j)})+\psi(\bm{\omega}_{2}^{(j)})\right)\right]^{2}\right),$ (11)
@@
where the expectation is taken over both the directions and lengths of the random walks. Suppose the directions remain independent, but the lengths $l_{1}^{(i)}$ and $l_{2}^{(i)}$ (and likewise $l_{1}^{(j)}$ and $l_{2}^{(j)}$) are to be coupled for variance reduction, analogously to the vector norms in Sec. 3. Let $\mathbb{E}_{\textrm{dirs}}$ denote the expectation over the walkers’ directions. We want to minimise:
$\mathbb{E}_{(l_{1}^{(i)},l_{2}^{(i)})\sim\mu,(l_{1}^{(j)},l_{2}^{(j)})\sim\mu}\mathbb{E}_{\textrm{dirs}}\left(\left[\left(\psi(\bm{\omega}_{1}^{(i)})+\psi(\bm{\omega}_{2}^{(i)})\right)^{\top}\left(\psi(\bm{\omega}_{1}^{(j)})+\psi(\bm{\omega}_{2}^{(j)})\right)\right]^{2}\right)\textrm{ for }{\mu\in\Lambda_{2}(\eta_{\textrm{G}})}.$ (12)